2 P(A) = = ,P(B) = 1,P(A1B) ,P(AB) + P(BA) Find P( AUB 3
Answer #1

P(8) PCA) 9 P(A)=1/4, PCB) = 1/7 , PLAIB)- P(ANB) - 7 P(ANB) P(BIA), PIANB) - 4 P(ANB) ::P(AIB)+P(B1A)= 343 = 7 PLAN B) +4P(A